OpenStudy (anonymous):

The quotient of a number and five is eight less than the number. What is the number? A. –2 B. 2 C. 3 D. 10

OpenStudy (e.mccormick):

Know what a quotient is?

OpenStudy (anonymous):

the result of division?

OpenStudy (e.mccormick):

Yes. So, if you have number / five = eight less than the number, how would you finish setting it up?

OpenStudy (anonymous):

2?

OpenStudy (e.mccormick):

No. Just take it one step at a time. Lets use n for number. So what is 8 less than n? Using +-/*, how can you write that?

OpenStudy (anonymous):

8-n

OpenStudy (e.mccormick):

That is n less than 8.

OpenStudy (anonymous):

so n-8

OpenStudy (e.mccormick):

Id I have 42 and I have 8 less than 42, then it is 42-8 = 36, so yes. n-8. Now put it back together with the n/5 and you get: \(\dfrac{n}{5}=n-8\)

OpenStudy (e.mccormick):

Can you solve it now?

OpenStudy (anonymous):

-2?

OpenStudy (anonymous):

or 10?

OpenStudy (e.mccormick):

\(\dfrac{n}{5}=n-8\) \(5\left( \dfrac{n}{5}\right) =5(n-8)\) n=5n-40 -4n=-40 n=10 Yah, I like that second answer a lot better!

OpenStudy (anonymous):

okay thank you!